OpenStudy (anonymous):

Which explicit formula was used to generate the first 5 terms of the sequence? 5, 3, 1, -1, -3

OpenStudy (cwrw238):

hint: each term (except the first) is obtained by adding -2 to the previous one.

OpenStudy (cwrw238):

are you familiar with sigma notation?

OpenStudy (anonymous):

um not to familiar

OpenStudy (anonymous):

a. an = 7n - 2 b. an = n + 5 c. an = -2n + 7 d. an = 2n - 7 im thinking its C

OpenStudy (cwrw238):

oh i see they want it in that form

OpenStudy (anonymous):

yes i forggot to mention

OpenStudy (cwrw238):

n is the cardinal number of the each number - that is first, second etc

OpenStudy (cwrw238):

an means the 'nth term'

OpenStudy (anonymous):

term number

OpenStudy (cwrw238):

which one of the options fits the first term 5 when n = 1?

OpenStudy (cwrw238):

yes or term number

OpenStudy (cwrw238):

clue - there are 2

OpenStudy (anonymous):

umm B i think

OpenStudy (cwrw238):

no if n = 1 , first term = 1 + 5 = 6

OpenStudy (cwrw238):

a and c give 5 when n = 1 but only one of these work for the other terms

OpenStudy (cwrw238):

so b and d are wrong

OpenStudy (anonymous):

oh! ok

OpenStudy (cwrw238):

follow?

OpenStudy (anonymous):

right thanks